OpenStudy (anonymous):

Can anyone derive dz/dt=(∂z/∂x)(dx/dt)+(∂z/∂y)(dy/dt). How can∂ cancel out with d. Aren't they not the same?

OpenStudy (anonymous):

Are you certain it isn't meant to be... \[\frac{ dz }{ dt }= \frac{ \delta z }{ \delta x } \frac{ \delta x }{ \delta t } + \frac{ \delta z }{ \delta y }\frac{ \delta y }{ \delta t }?\] that would make more sense as it would be the sums of partial differential equations..

OpenStudy (primeralph):

Someone tag me?

OpenStudy (primeralph):

@Dmitt That's a good question. I had the same question when I was in high school. Hold on.

OpenStudy (primeralph):

The proof I can generate deals with taking the impression of a curve on another. That's the way partials work. How good is your vector calculus?

OpenStudy (primeralph):

@Dmitt Can't stay her forever.